# Multimarked Spatial Search by Continuous-Time Quantum Walk

### 05 Jan 2024 v2

@@ 2 Multimarked framework
Let $G(V,E)$ be a finite connected graph with vertex set $V$ and edge set $E$. We associate $G$ with a Hilbert space $\mathcal{H}$ with computational basis $\{|v\rangle:v\in V\}$, as is usually done in the definition of the continuous-time quantum walk [1]. Let $W$ be the set of marked vertices. We consider the time-independent Hamiltonian of the form [3]
@@
where $A$ denotes the adjacency matrix of $G$, and $\gamma$ is a real and positive parameter. The spatial search algorithm starts with the initial state $\big{|}\psi(0)\big{\rangle}$, which is a normalized eigenvector of $-\gamma A$ associated with the largest eigenvalue of $A$. If $G$ is regular, $\big{|}\psi(0)\big{\rangle}$ is the uniform superposition of the computational basis
@@
The goal of the algorithm is to find the optimal values of parameters $t$ and $\gamma$ so that the success probability is as high as possible.
@@ 2.1 Eigenvalues and eigenvectors of the Hamiltonian
Given that $\sum_{\ell=0}^{k}P_{\ell}=I$, for a fixed marked vertex $w$, we obtain
$\big{\langle}w\big{|}\lambda\big{\rangle}=\sum_{\ell=0}^{k}\big{\langle}w\big{|}P_{\ell}\big{|}\lambda\big{\rangle}.$
Then, using Eq. (5), we derive
@@
where $\big{(}M_{ww^{\prime}}^{\lambda}\big{)}_{w,w^{\prime}\in W}$ is as defined in Eq. (3). Since $\lambda\not\in\sigma(-\gamma A)$, the vector $\big{\langle}w\big{|}\lambda\big{\rangle}\big{\rvert}_{w\in W}$ is nonzero by Proposition 1, and hence it is a 00-eigenvector of $M^{\lambda}$. It follows that
$\det\big{(}M^{\lambda}\big{)}=0.$
@@
which concludes the proof, considering Eq. (7). ∎
@@
Given that the initial state is $\big{|}\phi_{0}\big{\rangle}$, an eigenvector $\big{|}\lambda\big{\rangle}$ such that $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}=0$ does not affect the quantum walk dynamics. According to Eq. (7), the criterion to ensure that $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}\neq 0$ is
@@
Let $\lambda^{+}$ be the second smallest eigenvalue such that $\lambda^{+}\notin\sigma(-\gamma A)$ and besides there is an associated eigenvector $\big{|}\lambda^{+}\big{\rangle}$ satisfying $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}\neq 0$, that is, $\big{|}\lambda^{+}\big{\rangle}$ is obtained from an eigenvector of $M^{\lambda}$ that obeys condition (8). We claim that there exists a specific value of $\gamma$ for which the eigenvalue $-\gamma\phi_{0}$ of $-\gamma A$ is positioned midway between $\lambda^{-}$ and $\lambda^{+}$. Our approach uses $\lambda^{\pm}$ and this optimal $\gamma$ in the asymptotic limit. Essentially, for any graph $G(V,E)$ we claim the existence of a $\gamma$ value satisfying
@@
where $\epsilon>0$. Then, $\epsilon=|\lambda^{-}+\gamma\phi_{0}|$.
Before we present the proof of this claim, let us first offer some intuitive insights into its validity. Considering $H$ as defined in Eq. (1), when $\gamma=0$, the eigenvalues of $H$ are $-1$ (with multiplicity $|W|$) and 00 (with multiplicity $N-|W|$), where $N$ is the dimension of the Hilbert space. For small, nonzero $\gamma$, the eigenvalue of $-1$ splits into $|W|$ eigenvalues (which are not necessarily distinct) corresponding to different eigenvectors. Note that, based on a dimensionality analysis, at least one of these $|W|$ eigenvectors $\big{|}\lambda\big{\rangle}$ satisfies $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}\neq 0$, aligned with $\big{\langle}\lambda^{-}\big{|}\phi_{0}\big{\rangle}\neq 0$. In the case when $|W|\geq 2$, at least two eigenvalues will be different because $\lambda^{-}$, the smallest eigenvalue, is simple. Additionally, $\lambda^{+}$ is usually chosen to approach $-1$ as $\gamma$ decreases, automatically ensuring that $\lambda^{+}\notin\sigma(-\gamma A)$. However, in certain scenarios, $\lambda^{+}$ should approach 0, not $-1$, as $\gamma$ decreases, due to the constraint (8). In this case, we must ensure that $\lambda^{+}\notin\sigma(-\gamma A)$. Our claim remains valid when $\lambda^{+}$ approaches either $-1$ or 0 as $\gamma$ tends to 0.
When $\gamma$ is small, the eigenvalue $-\gamma\phi_{0}$ of $-\gamma A$ is near zero, placing it to the right of $\lambda^{+}$ on the real line spectrum or close to $\lambda^{+}$ in cases where the second eigenvalue is near zero when $\gamma$ goes to zero. Conversely, as $\gamma$ increases, $-\gamma\phi_{0}$ becomes increasingly aligned with $\lambda^{-}$, indicating an one-to-one correlation between the eigenvalues of $-\gamma A$ and $H$. We have established that for small $\gamma$, the eigenvalue $-\gamma\phi_{0}$ of $-\gamma A$ is positioned to the right or near the right of the range $[\lambda^{-},\lambda^{+}]$. As $\gamma$ increases, this eigenvalue shifts leftward, eventually nearing the left end of this range. Given that the determinants $|H-\lambda I|$ and $|-\gamma A-\phi I|$ are continuous real functions of $\gamma$, by applying the intermediate value theorem, we deduce that a specific $\gamma$ exists where $-\gamma\phi_{0}$ is exactly at the midpoint of $[\lambda^{-},\lambda^{+}]$.
@@
To illustrate the second eigenvalue $\lambda$ of $H$, which corresponds to an eigenvector $\big{|}\lambda\big{\rangle}$ satisfying $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}=0$, consider the complete bipartite graph $K_{n,n}$ with all marked vertices in one part. Here, the second smallest eigenvalue is $-1$, and each $(-1)$-eigenvector of $H$ derived from $K_{n,n}$ is orthogonal to $\big{|}\phi_{0}\big{\rangle}$. If $|W|\neq n$, then $\lambda^{+}$ is the third eigenvalue of $H$, approaching zero as $\gamma$ goes to zero. If $|W|=n$, the third eigenvalue of $H$ is in $\sigma(-\gamma A)$, making $\lambda^{+}$ the fourth and last eigenvalue of $H$.
@@ 2.2 Asymptotic analysis
In cases where one vertex is marked, several studies have employed a technique to analyze quantum-walk-based search algorithms. This technique uses two key eigenvectors of the evolution operator, ensuring significant overlap between both the initial and marked states and the space spanned by these eigenvectors [3, 4, 6]. In this work, we extend and generalize this technique to cases with multiple marked vertices. The asymptotic computational complexity of the search algorithm primarily depends on the eigenvalues $\lambda^{\pm}$, where $\lambda^{-}$ is the smallest eigenvalue of $H$, and $\lambda^{+}$ is the second smallest, satisfying the conditions $\lambda^{+}\notin\sigma(-\gamma A)$ and $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}\neq 0$. The optimal value of $\gamma$ is chosen according to the method outlined in Proposition 4, which positions $-\gamma\phi_{0}$ in the middle of the range $[\lambda^{-},\lambda^{+}]$. Since calculating this optimal $\gamma$ might be too difficult, we describe an alternative way to obtain $\gamma$, which asymptotically approaches the optimal one. Thus, we can write
@@
In this case, the gap between $\lambda^{-}$ and $\lambda^{+}$ is asymptotically $02\epsilon>$ 0. If the second smallest eigenvalue of $H$ naturally meets the restriction on $\lambda^{+}$, then $2\epsilon$ represents the spectral gap of $H$. We will detail a method to calculate $\gamma$ needed in Eq. (9) below.
@@
Proposition 4 ensures that Eq. (9) is applicable to any graph $G(V,E)$. Furthermore, this asymptotic analysis is valid for all graph classes where the gap between $\lambda^{-}$ and $\lambda^{+}$ approaches 0 as $N$ tends toward infinity. An example where the gap between $\lambda^{-}$ and $\lambda^{+}$ does not approach zero is the class of complete bipartite graphs $K_{n,n}$ with $n$ marked vertices in one part. In this case, using that $\phi_{0}=n$, the optimal parameters are $\lambda^{\pm}=(-1\pm\sqrt{2})/2$ and $\gamma=1/2n$. This yields a fixed $\epsilon=1/\sqrt{2}$. Note that the behavior of the gap depends on the placement of the marked vertices. In the same graph $K_{n,n}$, if the marked vertices are arranged in a different configuration, the gap approaches zero.
@@ 2.3 Calculation of the computational complexity
$p_{\text{exact}}(t)=\sum_{w\in W}\left|\sum_{\lambda}\mathrm{e}^{-\mathrm{i}\lambda t}\big{\langle}w\big{|}\lambda\big{\rangle}\big{\langle}\lambda\big{|}\psi(0)\big{\rangle}\right|^{2}.$
@@
$\big{|}\lambda^{-}\big{\rangle}$ is a normalized eigenvector of $H$ associated with the smallest root $\lambda^{-}$ of $\det(M^{\lambda})=0$ and $\big{|}\lambda^{+}\big{\rangle}$ with the second smallest root $\lambda^{+}$, subjected to the restrictions $\lambda^{+}\notin\sigma(-\gamma A)$ and $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}\neq 0$. Eigenvalue $\lambda^{-}$ is always simple but if eigenvalue $\lambda^{+}$ is not simple, which is uncommon, we must replace $\big{|}\lambda^{+}\big{\rangle}\big{\langle}\lambda^{+}\big{|}$ with the eigenprojection $P^{+}=\sum\big{|}\lambda^{+}\big{\rangle}\big{\langle}\lambda^{+}\big{|}$ in equation (11), where the sum runs over an orthonormal basis of the $\lambda^{+}$-eigenspace. The multiplicity of $\lambda^{+}$ is at most $|W|$.
The terms of $p_{\text{approx}}(t)$ are calculated as follows. The value of $\gamma$ is derived from the equation $a(\gamma)=0$ and $\epsilon$ from $\sqrt{b(\gamma)}$, as described in Sec. 2.2. Now, let us turn our attention to calculating $\big{|}\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\big{|}$. The approach for calculating $\big{|}\big{\langle}w\big{|}\lambda^{-}\big{\rangle}\big{|}$ follows a similar pattern. Remember that $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\bigr{\rvert}_{w\in W}$ is a 00-eigenvector of $M^{\lambda^{+}}$, with entries $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}$ for each $w\in W$. Given that $M^{\lambda^{+}}$ is already determined, we can identify a normalized 00-eigenvector of $M^{\lambda^{+}}$, which has entries denoted as $u(w)$ for each $w\in W$. In cases where $\lambda^{+}$ is not a simple eigenvalue, it becomes necessary to establish an orthonormal basis for the kernel of $M^{\lambda^{+}}$. Following this,
@@
$\frac{1}{\left|{c^{+}}\right|^{2}}=\sum_{\ell=0}^{k}\frac{1}{\left|\lambda^{+}+\gamma\phi_{\ell}\right|^{2}}\left\|\sum_{w\in W}u(w)P_{\ell}\big{|}w\big{\rangle}\right\|^{2}.$
@@
$\frac{1}{\left|{c^{+}}\right|}=\frac{1}{\epsilon}\left\|\sum_{w\in W}u(w)P_{0}\big{|}w\big{\rangle}\right\|+O(1).$
@@
where $\lambda^{\circ}$ represents the eigenvalue of $H$ immediately following $\lambda^{+}$. Using Eq. (5) and taking $\ell=0$, we find $\left(\lambda+\gamma\phi_{0}\right)\big{\langle}\psi(0)\big{|}\lambda\big{\rangle}=-\sum_{w}\big{\langle}\psi(0)\big{|}w\big{\rangle}\big{\langle}w\big{|}\lambda\big{\rangle}$. This leads to the upper bound
@@
Typically, the right-hand side tends to decrease. For example, in regular graphs, the summation on the right-hand side simplifies to $|W|^{2}/N$. If the gap between the eigenvalue $\lambda^{\circ}$ and $-\gamma\phi_{0}$ is sufficiently large, such that it diminishes more slowly than $1/\sqrt{N}$, then the overlap $|\big{\langle}\psi(0)\big{|}\lambda\big{\rangle}|^{2}$ for the excluded eigenvectors in Eq. (11) approaches zero. If the error approaches zero asymptotically, we can state that $p_{\text{exact}}(t)=p_{\text{approx}}(t)+o(1)$.
@@
$p_{\text{approx}}(t)=4\left|\big{\langle}\lambda^{+}\big{|}\psi(0)\big{\rangle}\right|^{2}\sum_{w\in W}\left|\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\right|^{2}\sin^{2}{\epsilon t}+o(1)+o(\epsilon t).$ (15)
@@
which shows that the gap between $\lambda^{\pm}$ determines the optimal running time. This expression leads to the success probability being
$p_{\mathrm{succ}}=4\left|\big{\langle}\lambda^{+}\big{|}\psi(0)\big{\rangle}\right|^{2}\sum_{w\in W}\left|\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\right|^{2}+o(1).$ (17)

### 17 Jan 2024 v3

@@ 2 Multimarked framework
Let $G(V,E)$ be a finite connected simple graph with vertex set $V$ and edge set $E$. We associate $G$ with a Hilbert space $\mathcal{H}$ with computational basis $\{|v\rangle:v\in V\}$, as is usually done in the definition of the continuous-time quantum walk [1]. Let $W$ be the set of marked vertices. We consider the time-independent Hamiltonian of the form [3]
@@
where $A$ denotes the adjacency matrix of $G(V,E)$, and $\gamma$ is a real and positive parameter. The spatial search algorithm starts with the initial state $\big{|}\psi(0)\big{\rangle}$, which is a normalized eigenvector of $-\gamma A$ associated with the largest eigenvalue of $A$. If $G$ is regular, $\big{|}\psi(0)\big{\rangle}$ is the uniform superposition of the computational basis
@@
Our goal is to find the optimal values of parameters $t$ and $\gamma$ so that the success probability is as high as possible.
@@ 2.1 Eigenvalues and eigenvectors of the Hamiltonian
Given that $\sum_{\ell=0}^{k}P_{\ell}=I$, for a fixed marked vertex $w$, we obtain $\big{\langle}w\big{|}\lambda\big{\rangle}=\sum_{\ell=0}^{k}\big{\langle}w\big{|}P_{\ell}\big{|}\lambda\big{\rangle}.$ Then, using Eq. (5), we derive
@@
where $\big{(}M_{ww^{\prime}}^{\lambda}\big{)}_{w,w^{\prime}\in W}$ is as defined in Eq. (3). Since $\lambda\not\in\sigma(-\gamma A)$, the vector $\big{\langle}w\big{|}\lambda\big{\rangle}\big{\rvert}_{w\in W}$ is nonzero by Proposition 1, and hence it is a 00-eigenvector of $M^{\lambda}$. It follows that $\det\big{(}M^{\lambda}\big{)}=0.$
@@
which concludes the proof, considering Eq. (7).
@@
Given that the initial state is $\big{|}\phi_{0}\big{\rangle}$, an eigenvector $\big{|}\lambda\big{\rangle}$ such that $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}=0$ does not affect the quantum walk dynamics. The criterion to ensure that $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}\neq 0$ when eigenvector $\big{|}\lambda\big{\rangle}$ is obtained using Eq. (7) is
@@
Let $\lambda^{+}$ denote the eigenvalue of $H$ that is the $(|W|+1)$-smallest, considering the multiplicity of each eigenvalue. If $\lambda^{+}$ is in the spectrum $\sigma(-\gamma A)$, or if the eigenvectors $\big{|}\lambda^{+}\big{\rangle}$ corresponding to $\lambda^{+}$ satisfy $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}=0$ for $\gamma>0$, meaning that $\big{|}\lambda^{+}\big{\rangle}$ is derived from a 00-eigenvector of $M^{\lambda}$ fulfilling condition (8), then we must select the next smallest eigenvalue. We claim that there exists a specific value of $\gamma$ for which the eigenvalue $-\gamma\phi_{0}$ of $-\gamma A$ is positioned midway between $\lambda^{-}$ and $\lambda^{+}$. Our approach uses $\lambda^{\pm}$ and this $\gamma$ in the asymptotic limit. Essentially, for any graph $G(V,E)$ we claim the existence of a $\gamma$ value satisfying
@@
where $|\lambda^{-}+\gamma\phi_{0}|=|\lambda^{+}+\gamma\phi_{0}|=\epsilon$.
Before we present the proof of this claim, let us first offer some intuitive insights into its validity. Considering $H$ as defined in Eq. (1), when $\gamma=0$, the eigenvalues of $H$ are $-1$ (with multiplicity $|W|$) and 00 (with multiplicity $N-|W|$), where $N$ is the dimension of the Hilbert space. For small, nonzero $\gamma$, the eigenvalue of $-1$ splits into $|W|$ eigenvalues (which are not necessarily distinct) corresponding to different eigenvectors. Note that, based on a dimensionality analysis, at least one of these $|W|$ eigenvectors $\big{|}\lambda\big{\rangle}$ satisfies $\big{\langle}\lambda\big{|}\phi_{0}\big{\rangle}\neq 0$, aligned with $\big{\langle}\lambda^{-}\big{|}\phi_{0}\big{\rangle}\neq 0$. In the case when $|W|\geq 2$, at least two eigenvalues will be different because $\lambda^{-}$, the smallest eigenvalue, is simple. As previously described, $\lambda^{+}$ represents the $(|W|+1)$-th eigenvalue of $H$, considering the multiplicity of each eigenvalue. Consequently, $\lambda^{+}$ is selected as the smallest eigenvalue of $H$ that approaches 0 when $\gamma$ tends to 0. Due to the stipulation in constraint (8) and the condition $\lambda^{+}\notin\sigma(-\gamma A)$, there are instances where $\lambda^{+}$ needs to be the $j$-th eigenvalue, where $j$ exceeds $|W|+1$. It is essential to avoid any eigenvalues that converge towards $-1$ as $\gamma$ decreases when choosing $\lambda^{+}$ because, in such scenarios, the gap between $\lambda^{-}$ and $\lambda^{+}$ becomes too small.
When $\gamma$ is small, the eigenvalue $-\gamma\phi_{0}$ of $-\gamma A$ is near zero, placing it close to $\lambda^{+}$ on the real line spectrum. Conversely, as $\gamma$ increases, $-\gamma\phi_{0}$ becomes increasingly aligned with $\lambda^{-}$, indicating an one-to-one correlation between the eigenvalues of $-\gamma A$ and $H$. We have established that for small $\gamma$, the eigenvalue $-\gamma\phi_{0}$ is positioned near the right of the range $[\lambda^{-},\lambda^{+}]$. As $\gamma$ increases, this eigenvalue shifts leftward, eventually nearing the left end of this range. Given that the determinants $|H-\lambda I|$ and $|-\gamma A-\phi I|$ are continuous real functions of $\gamma$, by applying the intermediate value theorem, we deduce that a specific $\gamma$ exists where $-\gamma\phi_{0}$ is exactly at the midpoint of $[\lambda^{-},\lambda^{+}]$.
@@
To illustrate an eigenvalue $\lambda$ to avoid because it is in $\sigma(-\gamma A)$, consider the complete bipartite graph $K_{n,n}$ with $n$ marked vertices in one part. The eigenvalues of $H$ are ordered as $\lambda^{-}<-1<0<\lambda^{+}$. The second smallest eigenvalue is $-1$ with a multiplicity of $n-1$. The next eigenvalue, 00, belongs to $\sigma(-\gamma A)$, making $\lambda^{+}$ the $(n+2)$-th and final eigenvalue of $H$.
@@ 2.2 Asymptotic analysis
In cases where one vertex is marked, several studies have employed a technique to analyze quantum-walk-based search algorithms. This technique uses two key eigenvectors of the evolution operator, ensuring significant overlap between both the initial and marked states and the space spanned by these eigenvectors [3, 4, 6]. In this work, we extend and generalize this technique to cases with multiple marked vertices. The asymptotic computational complexity of the search algorithm primarily depends on the eigenvalues $\lambda^{\pm}$, where $\lambda^{-}$ is the smallest eigenvalue of $H$, and $\lambda^{+}$ is the $(|W|+1)$-smallest, taking into account the multiplicity of each eigenvalue. Besides, $\lambda^{+}$ must satisfy the conditions $\lambda^{+}\notin\sigma(-\gamma A)$ and $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}\neq 0$ for $\gamma>0$. The value of $\gamma$ is described in Proposition 4, which positions $-\gamma\phi_{0}$ in the middle of the range $[\lambda^{-},\lambda^{+}]$. Since calculating this optimal $\gamma$ might be too difficult, we describe an alternative way to obtain $\gamma$, which asymptotically approaches the optimal one. Thus, we can write
@@
In this case, the gap between $\lambda^{-}$ and $\lambda^{+}$ is asymptotically $02\epsilon>$ 0. Note that $2\epsilon$ is larger than the spectral gap of $H$ unless $|W|=1$. We will detail a method to calculate $\gamma$ needed in Eq. (9) below.
@@
Proposition 4 ensures that Eq. (9) is applicable to any graph $G(V,E)$. Furthermore, this asymptotic analysis is valid for any graph such that $\epsilon\ll 1$ or for any graph class where the gap between $\lambda^{-}$ and $\lambda^{+}$ approaches 0 as $N$ tends toward infinity. An example where the gap between $\lambda^{-}$ and $\lambda^{+}$ does not approach zero is the class of complete bipartite graphs $K_{n,n}$ with $n$ marked vertices in one part. In this case, using that $\phi_{0}=n$, the optimal parameters are $\lambda^{\pm}=(-1\pm\sqrt{2})/2$ and $\gamma=1/2n$. This yields a fixed $\epsilon=1/\sqrt{2}$. Note that the behavior of the gap depends on the placement of the marked vertices. In the same graph $K_{n,n}$, if the marked vertices are arranged in a different configuration than the one above, the gap approaches zero.
@@ 2.3 Calculation of the computational complexity
$p_{\text{exact}}(t)=\sum_{w\in W}\Big{|}\sum_{\lambda}\mathrm{e}^{-\mathrm{i}\lambda t}\big{\langle}w\big{|}\lambda\big{\rangle}\big{\langle}\lambda\big{|}\psi(0)\big{\rangle}\Big{|}^{2}.$
@@
$\big{|}\lambda^{-}\big{\rangle}$ is a normalized eigenvector of $H$ associated with the smallest root $\lambda^{-}$ of $\det(M^{\lambda})=0$ and $\big{|}\lambda^{+}\big{\rangle}$ with the $(|W|+1)$-smallest root $\lambda^{+}$, subjected to the restrictions $\lambda^{+}\notin\sigma(-\gamma A)$ and $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}\neq 0$. Eigenvalue $\lambda^{-}$ is always simple but if eigenvalue $\lambda^{+}$ is not simple, which is uncommon, we must replace $\big{|}\lambda^{+}\big{\rangle}\big{\langle}\lambda^{+}\big{|}$ with the eigenprojection $P^{+}=\sum\big{|}\lambda^{+}\big{\rangle}\big{\langle}\lambda^{+}\big{|}$ in equation (11), where the sum runs over an orthonormal basis of the $\lambda^{+}$-eigenspace. The multiplicity of $\lambda^{+}$ is at most $|W|$.
Note that $p_{\text{exact}}(t)\geq p_{\text{approx}}(t)$. The terms of $p_{\text{approx}}(t)$ are calculated as follows. The value of $\gamma$ is derived from the equation $a(\gamma)=0$ and $\epsilon$ from $\sqrt{b(\gamma)}$, as described in Sec. 2.2. Now, let us turn our attention to calculating $\big{|}\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\big{|}$. The approach for calculating $\big{|}\big{\langle}w\big{|}\lambda^{-}\big{\rangle}\big{|}$ follows a similar pattern. Remember that $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}\bigr{\rvert}_{w\in W}$ is a 00-eigenvector of $M^{\lambda^{+}}$, with entries $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}$ for each $w\in W$. Given that $M^{\lambda^{+}}$ is already determined, we can identify a normalized 00-eigenvector of $M^{\lambda^{+}}$, which has entries denoted as $u(w)$ for each $w\in W$. In cases where $\lambda^{+}$ is not a simple eigenvalue, it becomes necessary to establish an orthonormal basis for the kernel of $M^{\lambda^{+}}$. Following this,
@@
$\frac{1}{\left|{c^{+}}\right|^{2}}=\sum_{\ell=0}^{k}\frac{1}{\left|\lambda^{+}+\gamma\phi_{\ell}\right|^{2}}\,\Big{\|}\sum_{w\in W}u(w)P_{\ell}\big{|}w\big{\rangle}\Big{\|}^{2}.$
@@
$\frac{1}{\left|{c^{+}}\right|}=\frac{1}{\epsilon}\,\Big{\|}\sum_{w\in W}u(w)P_{0}\big{|}w\big{\rangle}\Big{\|}+O(1).$
@@
where $\lambda^{\circ}$ represents the eigenvalue of $H$ immediately following $\lambda^{-}$. Using Eq. (5) and taking $\ell=0$, we find $\left(\lambda+\gamma\phi_{0}\right)\big{\langle}\psi(0)\big{|}\lambda\big{\rangle}=-\sum_{w}\big{\langle}\psi(0)\big{|}w\big{\rangle}\big{\langle}w\big{|}\lambda\big{\rangle}$. This leads to the upper bound
@@
Typically, the right-hand side tends to decrease. For example, in regular graphs, the summation on the right-hand side simplifies to $|W|^{2}/N$. If the gap between the eigenvalue $\lambda^{\circ}$ and $-\gamma\phi_{0}$ is sufficiently large, such that it diminishes more slowly than $1/\sqrt{N}$, then the overlap $|\big{\langle}\psi(0)\big{|}\lambda\big{\rangle}|^{2}$ for the excluded eigenvectors in Eq. (11) approaches zero. If the error approaches zero asymptotically, we can state that $p_{\text{exact}}(t)=p_{\text{approx}}(t)+o(1)$ and the framework described in this work is successful.
@@
$p_{\text{approx}}(t)=4\left|\big{\langle}\lambda^{-}\big{|}\psi(0)\big{\rangle}\right|^{2}\sum_{w\in W}\left|\big{\langle}w\big{|}\lambda^{-}\big{\rangle}\right|^{2}\sin^{2}{\epsilon t}+o(1)+o(\epsilon t).$ (15)
@@
which shows that the gap between $\lambda^{\pm}$ determines the optimal running time. A larger gap means a smaller runtime. This expression leads to the success probability being
$p_{\mathrm{succ}}=4\left|\big{\langle}\lambda^{-}\big{|}\psi(0)\big{\rangle}\right|^{2}\sum_{w\in W}\left|\big{\langle}w\big{|}\lambda^{-}\big{\rangle}\right|^{2}+o(1).$ (17)